# CM points on Shimura curves via QM-equivariant isogeny volcanoes

### v2

@@ 2.3.2. The field of moduli in the CM case
The convention used in González–Rotger is common in the literature, appearing in the work of Rotger and his collaborators and also in recent work of Padurariu–Schembri [PS22] in which the authors compute rational points on all Atkin–Lehner quotients of geometrically hyperelliptic Shimura curves. The difference in convention one takes is motivated by which moduli problem one chooses for the course moduli scheme $X_{0}^{D}(N)$: our choice of working with maximal orders results in having natural modular maps from $X_{0}^{D}(N)$ to $X_{0}^{D}(1)$ for all $N$, while working with Eichler orders of level $N$ naturally situates $X_{0}^{D}(N)$ as the base Shimura curve. Because we want to work with general level, we work with maximal orders. A main difference between our work and that of [GR06], beyond the generalization from squarefree $N$ to all positive integers $N$, is that we consider not just the CM points on a fixed curve $X_{0}^{D}(N)$ but the fiber of the covering $X_{0}^{D}(N)_{/\mathbb{Q}}\rightarrow X^{D}(1)_{/\mathbb{Q}}$ over any CM point.
@@ 3. QM-equivariant isogenies
In this section, we prove facts about QM-equivariant isogenies needed in the proceeding section. Much of what we do in both this section and the next is in strong analogy to the case of isogenies of elliptic curves over $\overline{\mathbb{Q}}$ studied in recent work of Clark and Clark–Saia [Cl23, CS23]. We provide proofs here for completeness and for clarity of said analogy.
@@
Let $F$ be a field of characteristic zero, and let $(A,\iota)$ be a QM abelian surface over $F$ which does not have CM by an order of discriminant $\Delta\in\{-3,-4\}$. For $\ell$ a prime number, the number of QM-cyclic $\ell$-isogenies with domain $(A,\iota)$ which are defined over $F$, up to isomorphism, is either $0,1,2,$ or $\ell+1$.
@@
Note that $\ell$ being prime means we are counting isomorphism classes of QM-cyclic $\ell$-isogenies. The hypotheses on $A$ are equivalent to $\textnormal{Aut}(A,\iota)=\{\pm 1\}.$ In this case, we have a bijective correspondence between isomorphism classes of QM-cyclic $\ell$-isogenies and non-trivial, proper cyclic $\mathcal{O}$-submodules of $A[\ell]$. Under this correspondence, the isogenies defined over $F$ correspond to $\text{Gal}(\overline{F}/F)$-stable submodules.
@@ 3.1. Compositions of QM-cyclic isogenies
The following result is in analogy with [Cl23, Prop. 3.2].
@@
The containment of part (1) is clear. The assumption that $(A_{2},\iota_{2})$ does not have $-3$ or $-4$ CM is equivalent to $\text{Aut}((A_{2},\iota_{2}))=\{\pm 1\}$, and in this case the reverse containment in part (2) follows by the same argument as in [Cl23, Prop. 3.2]. ∎
@@ 3.2. Reduction to prime power degrees
This follows as in the $D=1$ case of [Cl23, Prop. 3.5], using that $X_{0}^{D}(N)$ for $D>1$ is a cover of $X^{D}(1)$ with the same corresponding subgroup of $\text{GL}_{2}(\mathbb{Z}/N\mathbb{Z})/\{\pm 1\}$ as in the case of $X_{0}(N)\rightarrow X(1)$. ∎
@@ 4. QM-equivariant isogeny volcanoes
Fixing a prime $\ell$, we describe in this section CM components of $\ell$-isogeny graphs of QM abelian surfaces over $\overline{\mathbb{Q}}$. We will use the work of this section to study CM points on the curves $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$ for $a\in\mathbb{Z}^{+}$ and $D>1$, in analogy to the $D=1$ modular curve case of [Cl23, CS23].
This study, like that of [Cl23, CS23], is motivated by the foundational work on isogeny volcanoes over finite fields by Kohel in his PhD thesis [Koh96] and by Fouquet [Fou01] and Fouquet–Morain [FM02]. We also recommend, and will refer to, a more recent, expository account of isogeny volcanoes in the finite field setting by Sutherland [Sut13].
@@ 4.1. The isogeny graph of QM abelian surfaces
Fix a prime number $\ell$ and an imaginary quadratic field $K$. In [Cl23] and [CS23], the authors consider the multigraph with vertex set that of $j$-invariants of $K$-CM elliptic curves, and with edges corresponding to $\mathbb{C}$-isomorphism classes of cyclic $\ell$-isogenies.
@@
We now consider the corresponding isogeny of $K$-CM elliptic curves of conductor $f$ from (1). In doing so, [CS23, Lemma 4.1] tells us that we must have $\ell\nmid f$. There, the result is reached using the correspondence between horizontal $\ell$-isogenies of $\mathfrak{o}(f_{0})$-CM elliptic curves over $\mathbb{C}$ with proper $\mathfrak{o}(f_{0})$-ideals of norm $\ell$. This also gives us the count of horizontal isogenies mentioned; we have the count in the elliptic curve case as in [CS23], and from a horizontal isogeny of elliptic curves as in (1) we generate a QM-cyclic isogeny of our QM-abelian surfaces via the QM action. ∎
@@ 4.2. QM-equivariant isogeny volcanoes
For a prime number $\ell$, we define here the notion of an $\ell$-volcano. This notion for the most part agrees with that in the existing literature, with the only caveat being that in the original context of isogeny volcanoes over a finite field one has volcanoes of finite depth. In our case, working over an algebraically closed field as in [Cl23, CS23], we adjust the definition to allow for infinite depth volcanoes.
@@
This result is in exact analogy to [Cl23, Lemma 4.2], and the proof is as therein. ∎
@@ 5.1. Action of $\textnormal{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$
As a result, we have that the number of level $L^{\prime}$ vertices in a given orbit which are fixed by $\sigma$ is the same as the number of elements of $\textnormal{Pic}(\mathfrak{o}(\ell^{L^{\prime}}f_{0}))$ fixed by $\tau$. As shown in [Cl23, Prop. 2.6], this count is equal to #$\#\textnormal{Pic}(\mathfrak{o}(\ell^{L^{\prime}}f_{0}))[2]$, as $\tau$ acts on $\textnormal{Pic}(\mathfrak{o}(\ell^{L^{\prime}}f_{0}))$ by inverting ideals. ∎
@@ 5.2. The field of moduli of a QM-cyclic $\ell^{a}$ isogeny
We now explicitly analyze the Galois action in all cases as done in [Cl23, §5.3] and [CS23, §4.2] in the $D=1$ case. Borrowing the notation therein, for a specified $K,f_{0},$ and $\ell$ we let
@@ 5.3. Explicit description: $f_{0}^{2}\Delta_{K}<-4$
In the current section, we assume $f_{0}^{2}\Delta_{K}<-4$, such that each component of $\mathcal{G}^{D}_{K,\ell,f_{0}}$ has the structure of an $\ell$-volcano of infinite depth. This is in exact parallel to [Cl23, §5.3], baring the same structure of results.
@@
The number of descending edges from a given surface vertex is either $\ell+1$ or $\ell-1$ depending on whether $\ell$ is inert or split in $K$, hence is even in both cases. With our involution being of the form $\sigma=\tau\sigma_{0}$, a translated version of the argument of [CS23, Cor. 5.5] gives that each fixed surface vertex has at least one fixed descendant in level $1$. Therefore, each fixed surface vertex must have at least two fixed descendants in level $1$ by parity, giving the result. ∎
@@
Suppose to the contrary that $v_{L^{\prime}+1}$ and $w_{L^{\prime}+1}$ are $\sigma$-fixed neighbors of $v_{L^{\prime}}$ and $w_{L^{\prime}}$, respectively, in level $L^{\prime}+1$. We find that this cannot be the case as in [Cl23, Lemma 5.6 c]; this would imply that we have a QM-cyclic $2^{4}$-isogeny which , upon restriction, would provide a cyclic, real $2^{4}$-isogeny of elliptic curves with CM by $\Delta=2^{2L+2}f_{0}^{2}\Delta_{K}$. This in turn implies the existence of a primitive, proper real $\mathfrak{o}(2^{L+1}f_{0})$-ideal of index $16$, which does not exist. ∎
@@
This proof comes down to a simple argument about ideals of norm $2$ in $\mathfrak{o}(f_{0})$, as in [Cl23, Lemma 5.7]
@@ 5.4. Explicit description part II: $f_{0}^{2}\Delta_{K}\in\{-3,-4\}$
Keeping our notation from the previous section, we now assume $f_{0}=1$ and $\Delta_{K}\in\{-3,-4\}$. As mentioned earlier in this section, we always have $D(K)\neq 1$ in this case. Therefore, the action of $\text{Gal}(K(\ell^{L}f)/\mathbb{Q})$ on $V_{L}$ is free for all $L\geq 0$. This is splendid news for us; while the CM fields $\mathbb{Q}(i)$ and $\mathbb{Q}(\sqrt{-3})$ require extra attention at other points in this study, they cause absolutely no difficulties as far as determining the explicit Galois action on $\mathcal{G}^{D}_{K,\ell,1}$. This is to be compared with the $D=1$ case of [CS23, §4], wherein much care goes into defining and explicitly describing a meaningful action of complex conjugation on CM components of isogeny graphs in these cases.
Still, we provide here example figures of components of $\mathcal{G}_{K,\ell,1}$ (up to finite level $L$) for each case as reference for the reader for the path type analysis and enumeration done in §6. In these cases, edges from level 00 to $1$ have multiplicity as exposited in [CS23, §3] due to the presence of automorphisms that do not fix kernels of isogenies. We therefore do not have a one-to-one identification between edges and “dual” edges in this case, and so as in the referenced study we clearly denote edges with orientation and multiplicity between levels 00 and $1$.
@@ 6. CM points on $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$
If $D(K)=1$, then the path types showing up in our analysis of each $\mathcal{G}^{D}_{K,\ell,f_{0}}$ are exactly those appearing in [Cl23] and [CS23]. In this case, each graph $\mathcal{G}^{D}_{K,\ell,f_{0}}$ consists of $2^{b}$ copies of the analogous graph $\mathcal{G}_{K,\ell,f_{0}}$ from the $D=1$ modular curve case. Moreover, we have shown that the action of relevant involutions on each component is identical to the action of complex conjugation in the $D=1$ case, up to symmetry of our graphs. In each place where the isomorphism class of a residue field in the referenced $D=1$ analysis is a rational ring class field, we have in its place here some totally complex, index $2$ subfield of a ring class field as described in Theorem 2.8.
If at least one prime dividing $D$ is inert in $K$, i.e., if $D(K)>1$, then all of the residue fields of $\Delta$-CM points on $X^{D}(1)_{/\mathbb{Q}}$, and hence on $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$, are ring class fields. The path types showing up are exactly those in [CS23], but the counts will in general differ from the case of the previous paragraph. Specifically, a given path type in our analysis in the case of $D(K)=1$ consists of $m$ classes with corresponding residue field $K(f^{\prime})$ and $n$ classes with corresponding residue field an index $2$ subfield of $K(f^{\prime})$ for some $f^{\prime}\in\mathbb{Z}^{+}$ and $m,n\geq 0$. In the case of $D(K)>1$, the same path type then consists of $2m+n$ classes, each with corresponding residue field $K(f^{\prime})$.
@@
Suppose that $K=\mathbb{Q}(i)$ splits $B$, and consider the case of $\Delta=-4$ and $\ell^{a}=2$. We have $2^{b+1}$ closed-point equivalence classes of QM-cyclic $2$-isogenies of QM abelian surfaces with $-4$-CM. Each corresponding point on $X_{0}^{D}(2)_{/\mathbb{Q}}$ has residual degree $1$ over its image on $X^{D}(1)_{/\mathbb{Q}}$, having residue field $K$. Half of these classes, corresponding to self-loop edges at the surface, have no ramification, while each of the $2^{b}$ classes $C(\varphi)$ corresponding to a pair of descending edges to level $1$ has $e_{\varphi}=2$.
@@ 6.6. Primitive residue fields of CM points on $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$
The cases occurring here are in line with those in [Cl23] and [CS23], though here the primitive residue fields depend on whether $D(K)=1$. In particular, if some prime dividing $D$ is inert in $K$, then all residue fields of CM points on $X_{0}^{D}(\ell^{a})$ are ring class fields and hence there can only be one primitive residue field. This necessarily happens, for instance, if the class number of $K$ is odd. We provide Case 1.5b with the alternative title of “The dreaded case,” as in [Cl23], to warn the reader that it will have an important role in later results on primitive residue fields and degrees.

### v3

@@ 2.3.2. The field of moduli in the CM case
The convention used in González–Rotger is common in the literature, appearing in the work of Rotger and his collaborators and also in recent work of Padurariu–Schembri [PS23] in which the authors compute rational points on all Atkin–Lehner quotients of geometrically hyperelliptic Shimura curves. The difference in convention one takes is motivated by which moduli problem one chooses for the course moduli scheme $X_{0}^{D}(N)$: our choice of working with maximal orders results in having natural modular maps from $X_{0}^{D}(N)$ to $X_{0}^{D}(1)$ for all $N$, while working with Eichler orders of level $N$ naturally situates $X_{0}^{D}(N)$ as the base Shimura curve. Because we want to work with general level, we work with maximal orders. A main difference between our work and that of [GR06], beyond the generalization from squarefree $N$ to all positive integers $N$, is that we consider not just the CM points on a fixed curve $X_{0}^{D}(N)$ but the fiber of the covering $X_{0}^{D}(N)_{/\mathbb{Q}}\rightarrow X^{D}(1)_{/\mathbb{Q}}$ over any CM point.
@@ 3. QM-equivariant isogenies
In this section, we prove facts about QM-equivariant isogenies needed in the proceeding section. Much of what we do in both this section and the next is in strong analogy to the case of isogenies of elliptic curves over $\overline{\mathbb{Q}}$ studied in recent work of Clark and Clark–Saia [Cl22, CS22]. We provide proofs here for completeness and for clarity of said analogy.
@@
Let $F$ be a field of characteristic zero, and let $(A,\iota)$ be a QM abelian surface over $F$ which does not have CM by an order of discriminant $\Delta\in\{-3,-4\}$. For $\ell$ a prime number, the number of QM-cyclic $\ell$-isogenies with domain $(A,\iota)$ which are $\text{Gal}(\overline{F}/F)$-stable, up to isomorphism, is either $0,1,2,$ or $\ell+1$.
@@
Note that $\ell$ being prime means we are counting isomorphism classes of QM-cyclic $\ell$-isogenies. The hypotheses on $A$ are equivalent to $\textnormal{Aut}(A,\iota)=\{\pm 1\}.$ In this case, we have a bijective correspondence between isomorphism classes of QM-cyclic $\ell$-isogenies and non-trivial, proper cyclic $\mathcal{O}$-submodules of $A[\ell]$. Under this correspondence, the isogenies which are $\text{Gal}(\overline{F}/F)$-stable correspond to $\text{Gal}(\overline{F}/F)$-stable submodules.
@@ 3.1. Compositions of QM-cyclic isogenies
The following result is in analogy with [Cl22, Prop. 3.2].
@@
The containment of part (1) is clear. The assumption that $(A_{2},\iota_{2})$ does not have $-3$ or $-4$ CM is equivalent to $\text{Aut}((A_{2},\iota_{2}))=\{\pm 1\}$, and in this case the reverse containment in part (2) follows by the same argument as in [Cl22, Prop. 3.2]. ∎
@@ 3.2. Reduction to prime power degrees
This follows as in the $D=1$ case of [Cl22, Prop. 3.5], using that $X_{0}^{D}(N)$ for $D>1$ is a cover of $X^{D}(1)$ with the same corresponding subgroup of $\text{GL}_{2}(\mathbb{Z}/N\mathbb{Z})/\{\pm 1\}$ as in the case of $X_{0}(N)\rightarrow X(1)$. ∎
@@ 4. QM-equivariant isogeny volcanoes
Fixing a prime $\ell$, we describe in this section CM components of $\ell$-isogeny graphs of QM abelian surfaces over $\overline{\mathbb{Q}}$. We will use the work of this section to study CM points on the curves $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$ for $a\in\mathbb{Z}^{+}$ and $D>1$, in analogy to the $D=1$ modular curve case of [Cl22, CS22].
This study, like that of [Cl22, CS22], is motivated by the foundational work on isogeny volcanoes over finite fields by Kohel in his PhD thesis [Koh96] and by Fouquet [Fou01] and Fouquet–Morain [FM02]. We also recommend, and will refer to, a more recent, expository account of isogeny volcanoes in the finite field setting by Sutherland [Sut13].
@@ 4.1. The isogeny graph of QM abelian surfaces
Fix a prime number $\ell$ and an imaginary quadratic field $K$. In [Cl22] and [CS22], the authors consider the multigraph with vertex set that of $j$-invariants of $K$-CM elliptic curves, and with edges corresponding to $\mathbb{C}$-isomorphism classes of cyclic $\ell$-isogenies.
@@
We now consider the corresponding isogeny of $K$-CM elliptic curves of conductor $f$ from (1). In doing so, [CS22, Lemma 4.1] tells us that we must have $\ell\nmid f$. There, the result is reached using the correspondence between horizontal $\ell$-isogenies of $\mathfrak{o}(f_{0})$-CM elliptic curves over $\mathbb{C}$ with proper $\mathfrak{o}(f_{0})$-ideals of norm $\ell$. This also gives us the count of horizontal isogenies mentioned; we have the count in the elliptic curve case as in [CS22], and from a horizontal isogeny of elliptic curves as in (1) we generate a QM-cyclic isogeny of our QM-abelian surfaces via the QM action. ∎
@@ 4.2. QM-equivariant isogeny volcanoes
For a prime number $\ell$, we define here the notion of an $\ell$-volcano. This notion for the most part agrees with that in the existing literature, with the only caveat being that in the original context of isogeny volcanoes over a finite field one has volcanoes of finite depth. In our case, working over an algebraically closed field as in [Cl22, CS22], we adjust the definition to allow for infinite depth volcanoes.
@@
This result is in exact analogy to [Cl22, Lemma 4.2], and the proof is as therein. ∎
@@ 5.1. Action of $\textnormal{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$
As a result, we have that the number of level $L^{\prime}$ vertices in a given orbit which are fixed by $\sigma$ is the same as the number of elements of $\textnormal{Pic}(\mathfrak{o}(\ell^{L^{\prime}}f_{0}))$ fixed by $\tau$. As shown in [Cl22, Prop. 2.6], this count is equal to #$\#\textnormal{Pic}(\mathfrak{o}(\ell^{L^{\prime}}f_{0}))[2]$, as $\tau$ acts on $\textnormal{Pic}(\mathfrak{o}(\ell^{L^{\prime}}f_{0}))$ by inverting ideals. ∎
@@ 5.2. The field of moduli of a QM-cyclic $\ell^{a}$ isogeny
We now explicitly analyze the Galois action in all cases as done in [Cl22, §5.3] and [CS22, §4.2] in the $D=1$ case. Borrowing the notation therein, for a specified $K,f_{0},$ and $\ell$ we let
@@ 5.3. Explicit description: $f_{0}^{2}\Delta_{K}<-4$
In the current section, we assume $f_{0}^{2}\Delta_{K}<-4$, such that each component of $\mathcal{G}^{D}_{K,\ell,f_{0}}$ has the structure of an $\ell$-volcano of infinite depth. This is in exact parallel to [Cl22, §5.3], baring the same structure of results.
@@
The number of descending edges from a given surface vertex is either $\ell+1$ or $\ell-1$ depending on whether $\ell$ is inert or split in $K$, hence is even in both cases. With our involution being of the form $\sigma=\tau\sigma_{0}$, a translated version of the argument of [CS22, Cor. 5.5] gives that each fixed surface vertex has at least one fixed descendant in level $1$. Therefore, each fixed surface vertex must have at least two fixed descendants in level $1$ by parity, giving the result. ∎
@@
Suppose to the contrary that $v_{L^{\prime}+1}$ and $w_{L^{\prime}+1}$ are $\sigma$-fixed neighbors of $v_{L^{\prime}}$ and $w_{L^{\prime}}$, respectively, in level $L^{\prime}+1$. We find that this cannot be the case as in [Cl22, Lemma 5.6 c]; this would imply that we have a QM-cyclic $2^{4}$-isogeny which , upon restriction, would provide a cyclic, real $2^{4}$-isogeny of elliptic curves with CM by $\Delta=2^{2L+2}f_{0}^{2}\Delta_{K}$. This in turn implies the existence of a primitive, proper real $\mathfrak{o}(2^{L+1}f_{0})$-ideal of index $16$, which does not exist. ∎
@@
This proof comes down to a simple argument about ideals of norm $2$ in $\mathfrak{o}(f_{0})$, as in [Cl22, Lemma 5.7]
@@ 5.4. Explicit description part II: $f_{0}^{2}\Delta_{K}\in\{-3,-4\}$
Keeping our notation from the previous section, we now assume $f_{0}=1$ and $\Delta_{K}\in\{-3,-4\}$. As mentioned earlier in this section, we always have $D(K)\neq 1$ in this case. Therefore, the action of $\text{Gal}(K(\ell^{L}f)/\mathbb{Q})$ on $V_{L}$ is free for all $L\geq 0$. This is splendid news for us; while the CM fields $\mathbb{Q}(i)$ and $\mathbb{Q}(\sqrt{-3})$ require extra attention at other points in this study, they cause absolutely no difficulties as far as determining the explicit Galois action on $\mathcal{G}^{D}_{K,\ell,1}$. This is to be compared with the $D=1$ case of [CS22, §4], wherein much care goes into defining and explicitly describing a meaningful action of complex conjugation on CM components of isogeny graphs in these cases.
Still, we provide here example figures of components of $\mathcal{G}_{K,\ell,1}$ (up to finite level $L$) for each case as reference for the reader for the path type analysis and enumeration done in §6. In these cases, edges from level 00 to $1$ have multiplicity as exposited in [CS22, §3] due to the presence of automorphisms that do not fix kernels of isogenies. We therefore do not have a one-to-one identification between edges and “dual” edges in this case, and so as in the referenced study we clearly denote edges with orientation and multiplicity between levels 00 and $1$.
@@ 6. CM points on $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$
If $D(K)=1$, then the path types showing up in our analysis of each $\mathcal{G}^{D}_{K,\ell,f_{0}}$ are exactly those appearing in [Cl22] and [CS22]. In this case, each graph $\mathcal{G}^{D}_{K,\ell,f_{0}}$ consists of $2^{b}$ copies of the analogous graph $\mathcal{G}_{K,\ell,f_{0}}$ from the $D=1$ modular curve case. Moreover, we have shown that the action of relevant involutions on each component is identical to the action of complex conjugation in the $D=1$ case, up to symmetry of our graphs. In each place where the isomorphism class of a residue field in the referenced $D=1$ analysis is a rational ring class field, we have in its place here some totally complex, index $2$ subfield of a ring class field as described in Theorem 2.8.
If at least one prime dividing $D$ is inert in $K$, i.e., if $D(K)>1$, then all of the residue fields of $\Delta$-CM points on $X^{D}(1)_{/\mathbb{Q}}$, and hence on $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$, are ring class fields. The path types showing up are exactly those in [CS22], but the counts will in general differ from the case of the previous paragraph. Specifically, a given path type in our analysis in the case of $D(K)=1$ consists of $m$ classes with corresponding residue field $K(f^{\prime})$ and $n$ classes with corresponding residue field an index $2$ subfield of $K(f^{\prime})$ for some $f^{\prime}\in\mathbb{Z}^{+}$ and $m,n\geq 0$. In the case of $D(K)>1$, the same path type then consists of $2m+n$ classes, each with corresponding residue field $K(f^{\prime})$.
@@
Suppose that $K=\mathbb{Q}(i)$ splits $B$, and consider the case of $\Delta=-4$ and $\ell^{a}=2$. We have $2^{b+1}$ closed-point equivalence classes of QM-cyclic $2$-isogenies of QM abelian surfaces with $-4$-CM. Each corresponding point on $X_{0}^{D}(2)_{/\mathbb{Q}}$ has residue degree $1$ over its image on $X^{D}(1)_{/\mathbb{Q}}$, having residue field $K$. Half of these classes, corresponding to self-loop edges at the surface, have no ramification, while each of the $2^{b}$ classes $C(\varphi)$ corresponding to a pair of descending edges to level $1$ has $e_{\varphi}=2$.
@@ 6.6. Primitive residue fields of CM points on $X_{0}^{D}(\ell^{a})_{/\mathbb{Q}}$
The cases occurring here are in line with those in [Cl22] and [CS22], though here the primitive residue fields depend on whether $D(K)=1$. In particular, if some prime dividing $D$ is inert in $K$, then all residue fields of CM points on $X_{0}^{D}(\ell^{a})$ are ring class fields and hence there can only be one primitive residue field. This necessarily happens, for instance, if the class number of $K$ is odd. We provide Case 1.5b with the alternative title of “The dreaded case,” as in [Cl22], to warn the reader that it will have an important role in later results on primitive residue fields and degrees.